If I am 13 and my dad is 41 in how many years will he be exactly twice my age?
Answer by mananth(16946)

after x years you will be 1/2 dad's age
your age after x years
13+x
Dad's age after x years
41+x
13+x= 1/2(41+x)
2(13+x)=41+x
26+2x=41+x
2x-x=41-26
x=15
you will be 28 he will be 56